The population of the world in 2003 was 6.314 billion. It grew between 1975 and 2003 at an annual rate of 1.6%. Assume that it continues to grow at this rate.
a. Compute the doubling time.
b. Estimate the world population in 2048 and 2093 (assuming all other things remain unchanged).
1
Expert's answer
2019-10-14T08:53:51-0400

a. The doubling time using the rule of 70 is: t = 70/1.6 = 43.75 years.

Another way to calculate the doubling time is:

6.314×1.016^n = 12.628,

1.016^n = 2,

n = log1.016(2) = 43.67 years.

b. The world population in 2048 will be:

6.314×1.016^45 = 12.898 billion.

The world population in 2093 will be:

6.314×1.016^90 = 26.347 billion.